1. What mass of potassium permanganate must be dissolved to make 2.0L of a 4.0% (m/v) solution?
  2. Gasoline sold in Ontario must contain at least 5.0% (v/v) ethanol. How much ethanol is a driver likely to get when buying 30 L of gasoline?
  3. Calculate the molar concentrations of the ions in the solution. 20 g of calcium phosphate dissolved in 1.7 L of solution
  4. A solution contains 21.4 g of sodium nitrate, NaNO3(s), dissolved in 250 mL of solution. Find the percent (m/v) concentration of the solution.
  5. A solute has a solubility of 26g/100g H2O, if a solution was prepared to contain 115g/450g H2O, determine the nature of the solution.
  6. Calculate the molar concentration of a solution of 10 g of aluminum nitrite dissolved in 0.350 L of solution.
  7. Write the balanced net ionic equation for the following reactions (2)

Note: Hg2 has a charge of +2 in that it exists as Hg2+2

  1. Hg2(NO3)2(aq) + 2NaCl(aq) → Hg2Cl2(s) + 2NaNO3(aq)
  2. Pb(NO3)2(aq) + 2KCl(aq)  →2KNO3(aq)  +  PbCl2(s)   




1
Expert's answer
2021-04-19T04:34:56-0400

1 )

potassium permanganate mass = "2×\frac{4.0}{100}=0.08Kg"


2)

"(V/v)ethanol= 30×\frac{5}{100}=1.5kg"



3) Molar conc of calcium phosphate "=\frac{Moles}{Volume}"


"Moles = \frac{20}{310}=0.065 moles"


Molar concentrations "=\frac{0.065}{1.7}=0.038M"


4)

"Percent (m/v)=\frac{21.4}{250}×100=8.56" %